OpenStudy (anonymous):

The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers? A. 56 B. 41 C. 69 D. 55

OpenStudy (anonymous):

B.41 41+ (41+28) =110

OpenStudy (anonymous):

Let x and y be two numbers where y is the smaller of the two. x+y=110 x-y=28 <-- rearranged is x=28+y substitute x=28+y into the first equation (y+28)+y=110 2y+28=110 2y=82 y=41